2x over 5 + 1 over 3 = 7x -2 over 15?
Find the value of X

Answer by jndarrell(58)

Key step here is to find your common denominator of 3, 5, and 15. It's 15 bc all three of the denominators given can equally go into 15.
Next you will multiply your numerators by the same number it takes to make the denominator 15.
3(2x)/15 + 5(1)/15 = 7x-2/15
6x+5/15=7x-2/15
Since all terms are over 15, you can proceed to get x alone on one side.
Subtract 6x from both sides. It cancels the left side and leaves the right w/ 1x
Add 2 to both sides. It cancels out the right and leaves the left w/ 7
Your equation should now look like this:
x/15=7/15
To get rid of a denominator, you perform the opposite action which would be to multiply both sides by 15..resulting in cancelling both denominators
x=7
To check your work, plug the 7 back into the original equation for x.
:)